OpenStudy (anonymous):

5x-3y=3 2x-4y=-10 3x+y= ? a- -30 b- -13 c- -7 d-7 e-13

OpenStudy (wolf1728):

A) 5x-3y=3 B) 2x-4y=-10 Multiply equation B by -1 -2x + 4y =10 then add this to equation A 3x + y = 13 By doing this, we have found an equation that's exactly what we need. So, 3x + y = 13 and we didn't have to find the values of x and y.